What is the minimum value of f(x)=2x^2+5x-1
Answer by nerdybill(7384)

f(x)=2x^2+5x-1
minimun is when x is:
x = -b/(2a)
x = -5/(2*2)
x = -5/4
.
substitute back into:
f(-5/4)=2(-5/4)^2+5(-5/4)-1
f(-5/4)=2(25/16)+(-25/4)-1
f(-5/4)=(25/8)+(-25/4)-1
f(-5/4)=(25/8)+(-50/8)-8/8
f(-5/4)=-25/8-8/8
f(-5/4)=-33/8 (this is the minimum value)